OpenStudy (anonymous):

How to find the missing number in this series ..... , 30 , 20 , 35 , 23.33 , 38.38 , 25.56

OpenStudy (mathmath333):

\(\large\tt \color{black}{\dfrac{38.38}{25.56}=\dfrac{35}{23.33}=\dfrac{30}{20}=\dfrac{x}{30}=1.5}\) solve x \(\large\tt \color{black}{\dfrac{x}{30}=1.5}\)

OpenStudy (anonymous):

I did that before but the reason why I thought it was wrong because the fraction is equal to 1.50156495

OpenStudy (anonymous):

I mean the first fraction

OpenStudy (anonymous):

I could notice that as we are calculating the fractions we are approaching 1.5

OpenStudy (mathmath333):

1.50156495 0.00156495 -> this are not significant here

OpenStudy (anonymous):

So you think the missing number is 45?

OpenStudy (mathmath333):

not sure

OpenStudy (mathmath333):

45.00156495 ....lol might be

OpenStudy (anonymous):

They are very tricky